Suppose the demand for fast food workers can be defined a labor demand equation of LD = 95-3w. All fast food workers are covered by the minimum wage and it is binding. The minimum wage is increased from $7.25 to $9.00 per hour. Calculate the elasticity of demand for fast food workers over the relevant portion of the demand curve using the simple formula. Throughout your calculations round to two decimals.
Answer #1

LD at 7.25=95-3*7.25=95-21.75=73.25

LD at 9=95-3*9=95-27=68

% change in LD=(68-73.25)/73.25=-5.25/73.25=-7.17%

% change in minimum wage=(9-7.25)/7.25=1.75/7.25=24.14%

Elasticity =% change in LD/% changebin wage=-7.17%/24.14%=0.30

Hence elasticity=0.30